*Problem 15-5A X Your answer is incorrect. Try again. Phillips Corporations fiscal year ends on November 30. The following accounts are found in its job order cost accounting system for the first month of the new fiscal year. Other data 1. On December 1, two jobs were in process: Job No. 154 and Job No. 155. These jobs had combined direct materials costs of 2. During December, Job Nos. 156, 157, and 158 were started. On December 31, Job No. 158 was unfinished. This job had $9,950 and direct labor costs of $15,600. Overhead was applied at a rate that was 75% of direct labor cost. 5,300, plus manufacturing overhead. All jobs, except for Job No. 158, charges for direct materials $4,100 and direct labor were completed in December. On December 1, Job No. 153 was in the finished goods warehouse. It had a total cost of $5,500. On December 31, Job No. 157 was the only job finished that was not sold. It had a cost of $4,200. Manufacturing overhead was $1,220 underapplied in December. 3. 4.

Not sure what is missing, this is the entire question.
Answer #1
RAW MATERIAL
BEGINNING a 6590 REQUISITION 17250 a=17250+8165-18825
PURCHASES 18825 b=cost of direct labour+direct labour cost+OH(75%of labour)
ENDING BALANCE 8165 b=9950+15600+75%*15600
c=requisition of RM inventory-indirect RM
WIP c=17250-3400
BEGINNING b 37250 jobs completed f 53475 d=direct labour*75%
DM   c 13850 d=9000*75%
DL 9000 e=charges for dm+charges for dl+75%*dl (other data)
OVERHEAD d 6750 e=4100+5300+75%*5300
ENDING BALANCE e 13375 f=beginning WIP+DM+DL+OH-ENDING BALANCE OF WIP
f=37250+13850+9000+6750-13375
FINISHED GOODS INVENTORY g=5500 given in other data
BEGINNING g 5500 cogs 54775 h=see f above
completed jobs h 53475 i=beginning finished goods+job completed-ending finished goods
ENDING BALANCE j 4200 i=5500+53475-4200
j=4200 given in other data
factory labour k=factory labour incurred=12125
factory wages 12125 wage assigned k 12125 l=wage assigned-direct labour
l=12125-9000
manufacturing OH m=same as d
indirect material 3400 overhead applied d 6750
indirect labour l 3125
other overhead 1445